OpenStudy (anonymous):

I will medal someone What is the value of the expression? 2/3-(1/4)x(4/9) = A -5/9 B -1/6 C 1/6 D 5/9

OpenStudy (anonymous):

@whpalmer4

OpenStudy (faiqraees):

\[2/3 - 1/4 *4/9\] \[2/3 - 4/36\] \[2/3 - 1/9\] Divide both the numerator and denominator by 4 of the second fraction \[6/9 - 1/9\] Multiply both the numerator and denominator by 3 of the first fraction \[(6-1)/9\] \[5/9\]

OpenStudy (anonymous):

THANK YOU SO MUCH

OpenStudy (whpalmer4):

\[\frac{2}{3}-(\frac{1}{4})(\frac{4}{9}) = \frac{2}{3}-(\frac{1}{\cancel{4}})(\frac{\cancel{4}}{9})\]\[ = \frac{2}{3}-\frac{1}{9}=(\frac{3}{3})(\frac{2}{3})-\frac{1}{9}\]\[ = \frac{6}{9}-\frac{1}{9} = \frac{6-1}{9}\]\[=\frac{5}{9}\]